🔹BOILING POINT AND FREEZING POINT OF SOLUTION

1. a solution is made of 2.00g sucrose in 100.00 g H2O. (Molecular weight of sucrose is 342 g/mol). Calculate the (a) Boiling point and (b) freezing point of the solution 


🔸OSMOTIC PRESSURE 

1. What is the osmotic pressure of a solution made up of 2.00g sucrose (C12H22O11) in 100mL of solution at 30°C? 


1
Expert's answer
2021-05-12T06:30:17-0400

1.Sucrose mass = 2g


Molar mass of sucrose=342g/mol


No. of moles of sucrose "=\dfrac{2}{342}=0.0005847"


Mass of water= 100g

Molar mass of water = 18g/mol


Number of moles of water "=\dfrac{100}{18}=5.55 mol"


So using rault's law-


  "\dfrac{31.8-P_s}{31.8}=\dfrac{1.14}{32.94}"


"31.8-P_s=1.1005


\\[9pt]
 p_s=30.6995 mm hg"


2.

Osmotic pressure-


"\Pi= csT"


"=\dfrac{2}{100}\times 1000\times 0.0821\times 303
\\[9pt]
 =498.34 mm hg"